# ADS-Edit: A Multimodal Knowledge Editing Dataset for Autonomous Driving Systems

### v2

@@ 1. Introduction
The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024d; Wu et al., 2024a; Liu et al., 2024c; Team, 2024, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning. As a key foundation, LMMs have also found initial applications in Autonomous Driving Systems (ADS) (Xing et al., 2024; Sima et al., 2024; Huang et al., 2024a; Yao et al., 2024). However, the direct application of LMMs in ADS yields suboptimal results.
@@
To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to the model’s behavior, eschewing the substantial computational demands associated with full retraining. Unlike traditional fine-tuning, which risks catastrophic forgetting and demands extensive resources, Knowledge Editing facilitate rapid knowledge updates by selectively altering parameters associated with specific factual or contextual knowledge (Yao et al., 2023; Gupta et al., 2024; Zhang et al., 2024b; Fang et al., 2024a; Youssef et al., 2025; Jiang et al., 2025; Liu et al., 2024a; Yao et al., 2025). While preliminary explorations of Knowledge Editing in the area of LMMs have predominantly concentrated on the manipulation of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024d; Huang et al., 2024b), we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
@@
We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024d). Through our analysis, we obtained a series of interesting findings, including the universality of knowledge editing methods in updating knowledge across various scenarios, their ability to balance editing effectiveness and processing speed, particularly in the context of video data, and the remaining limitations in locality. Moreover, in real-world evaluation scenarios, it is evident that current editing techniques require further refinement and enhancement to meet the demands of autonomous driving systems.
@@ 4.1. Experimental settings
We conduct experiments on the most advanced LMMs to date, which are LLaVA-OneVision (Li et al., 2024d) and Qwen2-VL (Wang et al., 2024a). Both models utilize Qwen2-7B as their LLM component. Meanwhile, four classical knowledge editing baselines, including Prompt, AdaLora, GRACE, and WISE are employed to update the knowledge of LMMs. Our analysis covers both the Single Editing and Lifelong Editing settings to thoroughly assess the performance of different baselines in ADS-Edit. Single editing involves updating and evaluating the model immediately after receiving each individual driving data instance. In contrast, lifelong editing represents a more realistic driving scenario, where multiple driving data instances are collected during the vehicle’s moving, requiring continuous integration of this knowledge into the model followed by evaluation. We reasonably assume that the visual processing component (e.g., Vision Transformer) of LMMs provides reliable visual information. Therefore, we focus only on editing the LLM component to modify its understanding of visual input. To comprehensively evaluate the effectiveness of editing methods on ADS, we adopt two evaluation metrics: (1) Editing evaluation (Edit), where tokens are generated in a teacher-forcing manner and compared at the token level (Meng et al., 2022). (2) Real-world evaluation (Real), where the edited model performs free-form generation, and a LLM is prompted to assess the semantic consistency between the generated outputs and the ground-truth answers (Yang et al., 2025a). Due to cost and resource constraints, we report the Real metric only for the main single-edit experiment, while the remaining experiments are evaluated using the Edit metric by default.
@@ 5.1. Knowledge Editing
Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024b; Wu et al., 2024b; Xu et al., 2024; Li and Chu, 2024; Nie et al., 2024; Wei et al., 2025; Zhang et al., 2025; Ali et al., 2025; Hwang et al., 2025; Zhao et al., 2025; Markowitz et al., 2025; Yang et al., 2025b; Dong et al., 2025; Pan et al., 2025; Wang et al., 2025; Liu et al., [n. d.]; Zhang et al., 2024c; Xu et al., 2025; Bi et al., 2025; He et al., 2025; Li et al., 2025; Shen and Huang, 2025; Yang et al., 2025a). Following Yao et al. (2023), we systematize current methodologies into two principal paradigms: (1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024a; Tan et al., 2024; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023; Feng et al., 2025). (2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Jiang et al., 2024; Wang et al., 2024d; Hartvigsen et al., 2023; Li et al., 2025). The emergence of Large Multimodal Models (LMMs) has also advanced the field of multimodal knowledge editing. These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024; Pan et al., 2024; Wang et al., 2024c) or multimodal factual knowledge (Li et al., 2024a; Zhang et al., 2024d; Huang et al., 2024b). However, the application of multimodal knowledge editing in the context of domain-specific knowledge, like autonomous driving remains underexplored.
@@ 5.2. Large Multi-modal Models for Autonomous Driving
The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024; Cui et al., 2024) has spurred the development of advanced frameworks. CODA-LM (Li et al., 2024c) and DriveLM (Sima et al., 2024) integrate hierarchical reasoning and graph-structured visual QA frameworks. Meanwhile, LingoQA (Marcu et al., 2024) expands free-form QA capabilities by integrating action justification and scene understanding. However, these models still face challenges in real-world deployment due to static knowledge representations and modality imbalance.
@@ 6. Conclusion
Appendix A Metrics
Reliability.
Reliability is to evaluate the success of behavioral modification in the target driving scenarios. Intuitively, what we need is an updated with . To measure the reliability, we use the editing accuracy, as follows:
where refers to the parameters after editing.
Generality.
Given similar autonomous driving scenarios, generality assesses whether the edited model updates relevant knowledge within to retain the capacity for generalization, so that predict congruent outputs for equivalent inputs (e.g., rephrased textual quries, or similar autonomous driving’s videos). We evaluate text generality and multimodal generality as follows:
where presents the rephrased autonomous driving’s image or videos, refers to the rephrased textual queries, and denotes to in-scope objects of .
Locality.
After updating the model with domain-specific knowledge in autonomous driving, locality is used to evaluate whether knowledge editing preserves the model’s behavior on unrelated knowledge (e.g., unimodal factual knowledge and multimodal commonsense knowledge). Following (Cheng et al., 2023), we employ both text locality and multimodal locality to assess the stability of the editing process:
where and are distinct from . The tuples and represent samples drawn from and , respectively.
Appendix B Baselines
Prompt.
Directly alter the model’s behavior temporarily through prompts.
AdaLora.
LoRA introduces low-rank matrices into Transformer layers, fine-tuning only a small subset of parameters while keeping the majority of the model frozen. Building on LoRA, AdaLoRA (Zhang et al., 2023) improves LoRA by adaptively allocating the parameter budget based on the importance of weight matrices, using SVD to prune unimportant updates. In our work, we employ AdaLoRA as the baseline for efficient knowledge editing.
GRACE.
GRACE (Hartvigsen et al., 2023) introduces a discrete key-value codebook approach, where edits are cached as latent space mappings without altering the original model weights. This method enables thousands of sequential edits by storing corrections in a retrieval-based codebook, ensuring minimal interference with unrelated inputs and strong locality . However, GRACE’s reliance on non-parametric representations limits its ability to generalize beyond memorized edits, as it struggles to integrate new knowledge into the model’s reasoning process.
WISE.
WISE (Wang et al., 2024d) uses a dual-memory architecture comprising a main memory for pretrained knowledge and a side memory for editing. A routing mechanism decides which memory to use during inference. Additionally, knowledge is partitioned into multiple subspaces through sharding to facilitate conflict-free edits, with subsequent merging of the shards into a unified side memory.
Appendix C Lifelong Editing Results
In real-world autonomous driving scenarios, sequential knowledge updates are often required. We sequentially test the effects of lifelong editing 1, 250, 500, 750, and 1000 times across four baseline methods: Prompt, AdaLora, GRACE, and WISE. The lifelong editing results of Qwen2-VL and LLaVA-Onevision are shown in Figure 7 and Figure 8, respectively. Regarding Reliability and Generality metrics, the performance of the four baselines shows a gradual decline, albeit with some fluctuations. WISE, as a memory-augmented approach, effectively alleviates the knowledge forgetting phenomenon observed in AdaLora, which relies on parameter updates. Although the Prompt method demonstrates also effective lifelong editing performance at 250 and 500 times updates, it leds to an Out-of-Memory (OOM) error after 750 times updates. Due to the long tokens of multimodal inputs, the codebook struggles to differentiate between distinct representation information, resulting in suboptimal lifelong editing performance for GRACE. Methods that preserve the original parameters, such as WISE and GRACE, achieved 100% performance on the Locality metric. Due to interference from additional inputs, the Prompt method performs poorly on the Locality metric, with even worse results observed in multimodal locality. AdaLoRA employs low-rank matrix-based parameter updates, which minimally affect the original parameters, resulting in a strong performance on the locality metric.
Appendix D Benchmark Construction Details
Appendix E Case of ADS-Edit benchmark

### v3

@@ 1. Introduction
The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024d; Team, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning. As a key foundation, LMMs have also found initial applications in Autonomous Driving Systems (ADS) (Xing et al., 2024; Sima et al., 2024; Huang et al., 2024a; Yao et al., 2024). However, the direct application of LMMs in ADS yields suboptimal results.
@@
To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to the model’s behavior, eschewing the substantial computational demands associated with full retraining. Unlike traditional fine-tuning, which risks catastrophic forgetting and demands extensive resources, Knowledge Editing facilitate rapid knowledge updates by selectively altering parameters associated with specific factual or contextual knowledge (Yao et al., 2023; Gupta et al., 2024; Zhang et al., 2024b; Fang et al., 2024a; Youssef et al., 2025; Jiang et al., 2025; Liu et al., 2024a; Yao et al., 2025). While preliminary explorations of Knowledge Editing in the area of LMMs have predominantly concentrated on the manipulation of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024c; Huang et al., 2024b), we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
@@
We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024c). Through our analysis, we obtained a series of interesting findings, including the universality of knowledge editing methods in updating knowledge across various scenarios, their ability to balance editing effectiveness and processing speed, particularly in the context of video data, and the remaining limitations in locality. Moreover, in real-world evaluation scenarios, it is evident that current editing techniques require further refinement and enhancement to meet the demands of autonomous driving systems.
@@ 4.1. Experimental settings
We conduct experiments on the most advanced LMMs to date, which are LLaVA-OneVision (Li et al., 2024d) and Qwen2-VL (Wang et al., 2024a). Both models utilize Qwen2-7B as their LLM component. Meanwhile, four classical knowledge editing baselines, including Prompt, AdaLora, GRACE, and WISE are employed to update the knowledge of LMMs. Our analysis covers both the Single Editing and Lifelong Editing settings to thoroughly assess the performance of different baselines in ADS-Edit. Single editing involves updating and evaluating the model immediately after receiving each individual driving data instance. In contrast, lifelong editing represents a more realistic driving scenario, where multiple driving data instances are collected during the vehicle’s moving, requiring continuous integration of this knowledge into the model followed by evaluation. We reasonably assume that the visual processing component (e.g., Vision Transformer) of LMMs provides reliable visual information. Therefore, we focus only on editing the LLM component to modify its understanding of visual input. To comprehensively evaluate the effectiveness of editing methods on ADS, we adopt two evaluation metrics: (1) Editing evaluation (Edit), where tokens are generated in a teacher-forcing manner and compared at the token level (Meng et al., 2022). (2) Real-world evaluation (Real), where the edited model performs free-form generation, and a LLM is prompted to assess the semantic consistency between the generated outputs and the ground-truth answers (Yang et al., 2025). Due to cost and resource constraints, we report the Real metric only for the main single-edit experiment, while the remaining experiments are evaluated using the Edit metric by default.
@@ 5.1. Knowledge Editing
Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024b; Wu et al., 2024; Xu et al., 2024). Following Yao et al. (2023), we systematize current methodologies into two principal paradigms: (1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024a; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023). (2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Wang et al., 2024c; Hartvigsen et al., 2023). The emergence of Large Multimodal Models (LMMs) has also advanced the field of multimodal knowledge editing. These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023) or multimodal factual knowledge (Li et al., 2024a; Huang et al., 2024b). However, the application of multimodal knowledge editing in the context of domain-specific knowledge, like autonomous driving remains underexplored.
@@ 5.2. Large Multi-modal Models for Autonomous Driving
The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024) has spurred the development of advanced frameworks. CODA-LM (Li et al., 2024c) and DriveLM (Sima et al., 2024) integrate hierarchical reasoning and graph-structured visual QA frameworks. Meanwhile, LingoQA (Marcu et al., 2024) expands free-form QA capabilities by integrating action justification and scene understanding. However, these models still face challenges in real-world deployment due to static knowledge representations and modality imbalance.
@@ 6. Conclusion
Acknowledgements.
We would like to express our sincere gratitude to the anonymous reviewers for their thoughtful and constructive feedback. This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023- 00138), Ningbo Natural Science Foundation (2024J020), Yongjiang Talent Introduction Programme (2021A-156-G), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University. This work was supported by Ant Group and Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph.References
- (1)
- Antol et al. (2015) Stanislaw Antol, Aishwarya Agrawal, Jiasen Lu, Margaret Mitchell, Dhruv Batra, C. Lawrence Zitnick, and Devi Parikh. 2015. VQA: Visual Question Answering. In International Conference on Computer Vision (ICCV).
- Cai and Cao (2024) Yuchen Cai and Ding Cao. 2024. O-Edit: Orthogonal Subspace Editing for Language Model Sequential Editing. CoRR abs/2410.11469 (2024). doi:10.48550/ARXIV.2410.11469 arXiv:2410.11469
- Chen (2024) Huajun Chen. 2024. Large Knowledge Model: Perspectives and Challenges. Data Intelligence 3 (2024).
- Cheng et al. (2023) Siyuan Cheng, Bozhong Tian, Qingbin Liu, Xi Chen, Yongheng Wang, Huajun Chen, and Ningyu Zhang. 2023. Can We Edit Multimodal Large Language Models?. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 13877–13888. doi:10.18653/V1/2023.EMNLP-MAIN.856
- Dettmers et al. (2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. 2023. QLoRA: Efficient Finetuning of Quantized LLMs. In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023, NeurIPS 2023, New Orleans, LA, USA, December 10 - 16, 2023, Alice Oh, Tristan Naumann, Amir Globerson, Kate Saenko, Moritz Hardt, and Sergey Levine (Eds.). http://papers.nips.cc/paper_files/paper/2023/hash/1feb87871436031bdc0f2beaa62a049b-Abstract-Conference.html
- et al (2024) DeepSeek-AI et al. 2024. DeepSeek-V3 Technical Report. arXiv:2412.19437 [cs.CL] https://arxiv.org/abs/2412.19437
- Fang et al. (2024a) Junfeng Fang, Houcheng Jiang, Kun Wang, Yunshan Ma, Xiang Wang, Xiangnan He, and Tat-Seng Chua. 2024a. AlphaEdit: Null-Space Constrained Knowledge Editing for Language Models. CoRR abs/2410.02355 (2024). doi:10.48550/ARXIV.2410.02355 arXiv:2410.02355
- Fang et al. (2024b) Junfeng Fang, Houcheng Jiang, Kun Wang, Yunshan Ma, Xiang Wang, Xiangnan He, and Tat-Seng Chua. 2024b. AlphaEdit: Null-Space Constrained Knowledge Editing for Language Models. CoRR abs/2410.02355 (2024). doi:10.48550/ARXIV.2410.02355 arXiv:2410.02355
- Gupta et al. (2024) Akshat Gupta, Anurag Rao, and Gopala Anumanchipalli. 2024. Model editing at scale leads to gradual and catastrophic forgetting. arXiv preprint arXiv:2401.07453 (2024).
- Hartvigsen et al. (2023) Tom Hartvigsen, Swami Sankaranarayanan, Hamid Palangi, Yoon Kim, and Marzyeh Ghassemi. 2023. Aging with GRACE: Lifelong Model Editing with Discrete Key-Value Adaptors. In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023, NeurIPS 2023, New Orleans, LA, USA, December 10 - 16, 2023, Alice Oh, Tristan Naumann, Amir Globerson, Kate Saenko, Moritz Hardt, and Sergey Levine (Eds.). http://papers.nips.cc/paper_files/paper/2023/hash/95b6e2ff961580e03c0a662a63a71812-Abstract-Conference.html
- Hu et al. (2022) Edward J. Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2022. LoRA: Low-Rank Adaptation of Large Language Models. In The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022. OpenReview.net. https://openreview.net/forum?id=nZeVKeeFYf9
- Huang et al. (2024b) Han Huang, Haitian Zhong, Tao Yu, Qiang Liu, Shu Wu, Liang Wang, and Tieniu Tan. 2024b. VLKEB: A Large Vision-Language Model Knowledge Editing Benchmark. In Advances in Neural Information Processing Systems 38: Annual Conference on Neural Information Processing Systems 2024, NeurIPS 2024, Vancouver, BC, Canada, December 10 - 15, 2024, Amir Globersons, Lester Mackey, Danielle Belgrave, Angela Fan, Ulrich Paquet, Jakub M. Tomczak, and Cheng Zhang (Eds.). http://papers.nips.cc/paper_files/paper/2024/hash/1198b53fa686831d5f0c0860d7ec4f34-Abstract-Datasets_and_Benchmarks_Track.html
- Huang et al. (2024a) Zhijian Huang, Chengjian Feng, Feng Yan, Baihui Xiao, Zequn Jie, Yujie Zhong, Xiaodan Liang, and Lin Ma. 2024a. DriveMM: All-in-One Large Multimodal Model for Autonomous Driving. CoRR abs/2412.07689 (2024). doi:10.48550/ARXIV.2412.07689 arXiv:2412.07689
- Jiang et al. (2025) Houcheng Jiang, Junfeng Fang, Ningyu Zhang, Guojun Ma, Mingyang Wan, Xiang Wang, Xiangnan He, and Tat-seng Chua. 2025. AnyEdit: Edit Any Knowledge Encoded in Language Models. arXiv preprint arXiv:2502.05628 (2025).
- Kwiatkowski et al. (2019) Tom Kwiatkowski, Jennimaria Palomaki, Olivia Redfield, Michael Collins, Ankur P. Parikh, Chris Alberti, Danielle Epstein, Illia Polosukhin, Jacob Devlin, Kenton Lee, Kristina Toutanova, Llion Jones, Matthew Kelcey, Ming-Wei Chang, Andrew M. Dai, Jakob Uszkoreit, Quoc Le, and Slav Petrov. 2019. Natural Questions: a Benchmark for Question Answering Research. Trans. Assoc. Comput. Linguistics 7 (2019), 452–466. doi:10.1162/TACL_A_00276
- Li et al. (2024d) Bo Li, Yuanhan Zhang, Dong Guo, Renrui Zhang, Feng Li, Hao Zhang, Kaichen Zhang, Yanwei Li, Ziwei Liu, and Chunyuan Li. 2024d. LLaVA-OneVision: Easy Visual Task Transfer. CoRR abs/2408.03326 (2024). doi:10.48550/ARXIV.2408.03326 arXiv:2408.03326
- Li et al. (2024a) Jiaqi Li, Miaozeng Du, Chuanyi Zhang, Yongrui Chen, Nan Hu, Guilin Qi, Haiyun Jiang, Siyuan Cheng, and Bozhong Tian. 2024a. MIKE: A New Benchmark for Fine-grained Multimodal Entity Knowledge Editing. In Findings of the Association for Computational Linguistics, ACL 2024, Bangkok, Thailand and virtual meeting, August 11-16, 2024, Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, 5018–5029. doi:10.18653/V1/2024.FINDINGS-ACL.298
- Li et al. (2024b) Yanze Li, Wenhua Zhang, Kai Chen, Yanxin Liu, Pengxiang Li, Ruiyuan Gao, Lanqing Hong, Meng Tian, Xinhai Zhao, Zhenguo Li, et al. 2024b. Automated Evaluation of Large Vision-Language Models on Self-driving Corner Cases. arXiv preprint arXiv:2404.10595 (2024).
- Li et al. (2024c) Yanze Li, Wenhua Zhang, Kai Chen, Yanxin Liu, Pengxiang Li, Ruiyuan Gao, Lanqing Hong, Meng Tian, Xinhai Zhao, Zhenguo Li, Dit-Yan Yeung, Huchuan Lu, and Xu Jia. 2024c. Automated Evaluation of Large Vision-Language Models on Self-driving Corner Cases. CoRR abs/2404.10595 (2024). doi:10.48550/ARXIV.2404.10595 arXiv:2404.10595
- Liu et al. (2024a) Zeyu Leo Liu, Shrey Pandit, Xi Ye, Eunsol Choi, and Greg Durrett. 2024a. Codeupdatearena: Benchmarking knowledge editing on api updates. arXiv preprint arXiv:2407.06249 (2024).
- Liu et al. (2024b) Zeyu Leo Liu, Shrey Pandit, Xi Ye, Eunsol Choi, and Greg Durrett. 2024b. CodeUpdateArena: Benchmarking Knowledge Editing on API Updates. CoRR abs/2407.06249 (2024). doi:10.48550/ARXIV.2407.06249 arXiv:2407.06249
- Ma et al. (2024) Yaohui Ma, Xiaopeng Hong, Shizhou Zhang, Huiyun Li, Zhilin Zhu, Wei Luo, and Zhiheng Ma. 2024. ComprehendEdit: A Comprehensive Dataset and Evaluation Framework for Multimodal Knowledge Editing. CoRR abs/2412.12821 (2024). doi:10.48550/ARXIV.2412.12821 arXiv:2412.12821
- Marcu et al. (2024) Ana-Maria Marcu, Long Chen, Jan Hünermann, Alice Karnsund, Benoît Hanotte, Prajwal Chidananda, Saurabh Nair, Vijay Badrinarayanan, Alex Kendall, Jamie Shotton, Elahe Arani, and Oleg Sinavski. 2024. LingoQA: Visual Question Answering for Autonomous Driving. In Computer Vision - ECCV 2024 - 18th European Conference, Milan, Italy, September 29-October 4, 2024, Proceedings, Part LXXVII (Lecture Notes in Computer Science, Vol. 15135), Ales Leonardis, Elisa Ricci, Stefan Roth, Olga Russakovsky, Torsten Sattler, and Gül Varol (Eds.). Springer, 252–269. doi:10.1007/978-3-031-72980-5_15
- Meng et al. (2022) Kevin Meng, David Bau, Alex Andonian, and Yonatan Belinkov. 2022. Locating and Editing Factual Associations in GPT. In Advances in Neural Information Processing Systems 35: Annual Conference on Neural Information Processing Systems 2022, NeurIPS 2022, New Orleans, LA, USA, November 28 - December 9, 2022, Sanmi Koyejo, S. Mohamed, A. Agarwal, Danielle Belgrave, K. Cho, and A. Oh (Eds.). http://papers.nips.cc/paper_files/paper/2022/hash/6f1d43d5a82a37e89b0665b33bf3a182-Abstract-Conference.html
- Meng et al. (2023) Kevin Meng, Arnab Sen Sharma, Alex J. Andonian, Yonatan Belinkov, and David Bau. 2023. Mass-Editing Memory in a Transformer. In The Eleventh International Conference on Learning Representations, ICLR 2023, Kigali, Rwanda, May 1-5, 2023. OpenReview.net. https://openreview.net/forum?id=MkbcAHIYgyS
- Mitchell et al. (2022) Eric Mitchell, Charles Lin, Antoine Bosselut, Chelsea Finn, and Christopher D. Manning. 2022. Fast Model Editing at Scale. In The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022. OpenReview.net. https://openreview.net/forum?id=0DcZxeWfOPt
- Qian et al. (2024) Tianwen Qian, Jingjing Chen, Linhai Zhuo, Yang Jiao, and Yu-Gang Jiang. 2024. NuScenes-QA: A Multi-Modal Visual Question Answering Benchmark for Autonomous Driving Scenario. In Thirty-Eighth AAAI Conference on Artificial Intelligence, AAAI 2024, Thirty-Sixth Conference on Innovative Applications of Artificial Intelligence, IAAI 2024, Fourteenth Symposium on Educational Advances in Artificial Intelligence, EAAI 2014, February 20-27, 2024, Vancouver, Canada, Michael J. Wooldridge, Jennifer G. Dy, and Sriraam Natarajan (Eds.). AAAI Press, 4542–4550. doi:10.1609/AAAI.V38I5.28253
- Sima et al. (2024) Chonghao Sima, Katrin Renz, Kashyap Chitta, Li Chen, Hanxue Zhang, Chengen Xie, Jens Beißwenger, Ping Luo, Andreas Geiger, and Hongyang Li. 2024. DriveLM: Driving with Graph Visual Question Answering. In Computer Vision - ECCV 2024 - 18th European Conference, Milan, Italy, September 29-October 4, 2024, Proceedings, Part LII (Lecture Notes in Computer Science, Vol. 15110), Ales Leonardis, Elisa Ricci, Stefan Roth, Olga Russakovsky, Torsten Sattler, and Gül Varol (Eds.). Springer, 256–274. doi:10.1007/978-3-031-72943-0_15
- Team (2025) Qwen Team. 2025. Qwen2.5-VL. https://qwenlm.github.io/blog/qwen2.5-vl/
- Wang et al. (2024a) Peng Wang, Shuai Bai, Sinan Tan, Shijie Wang, Zhihao Fan, Jinze Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, Yang Fan, Kai Dang, Mengfei Du, Xuancheng Ren, Rui Men, Dayiheng Liu, Chang Zhou, Jingren Zhou, and Junyang Lin. 2024a. Qwen2-VL: Enhancing Vision-Language Model’s Perception of the World at Any Resolution. CoRR abs/2409.12191 (2024). doi:10.48550/ARXIV.2409.12191 arXiv:2409.12191
- Wang et al. (2024c) Peng Wang, Zexi Li, Ningyu Zhang, Ziwen Xu, Yunzhi Yao, Yong Jiang, Pengjun Xie, Fei Huang, and Huajun Chen. 2024c. WISE: Rethinking the Knowledge Memory for Lifelong Model Editing of Large Language Models. In The Thirty-eighth Annual Conference on Neural Information Processing Systems. https://openreview.net/forum?id=VJMYOfJVC2
- Wang et al. (2024b) Yiwei Wang, Muhao Chen, Nanyun Peng, and Kai-Wei Chang. 2024b. DeepEdit: Knowledge Editing as Decoding with Constraints. CoRR abs/2401.10471 (2024). doi:10.48550/ARXIV.2401.10471 arXiv:2401.10471
- Wu et al. (2024) Xiaobao Wu, Liangming Pan, William Yang Wang, and Anh Tuan Luu. 2024. AKEW: Assessing Knowledge Editing in the Wild. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, EMNLP 2024, Miami, FL, USA, November 12-16, 2024, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, 15118–15133. https://aclanthology.org/2024.emnlp-main.843
- Xing et al. (2024) Shuo Xing, Chengyuan Qian, Yuping Wang, Hongyuan Hua, Kexin Tian, Yang Zhou, and Zhengzhong Tu. 2024. OpenEMMA: Open-Source Multimodal Model for End-to-End Autonomous Driving. CoRR abs/2412.15208 (2024). doi:10.48550/ARXIV.2412.15208 arXiv:2412.15208
- Xu et al. (2024) Derong Xu, Ziheng Zhang, Zhihong Zhu, Zhenxi Lin, Qidong Liu, Xian Wu, Tong Xu, Wanyu Wang, Yuyang Ye, Xiangyu Zhao, Enhong Chen, and Yefeng Zheng. 2024. Editing Factual Knowledge and Explanatory Ability of Medical Large Language Models. In Proceedings of the 33rd ACM International Conference on Information and Knowledge Management, CIKM 2024, Boise, ID, USA, October 21-25, 2024, Edoardo Serra and Francesca Spezzano (Eds.). ACM, 2660–2670. doi:10.1145/3627673.3679673
- Yang et al. (2025) Wanli Yang, Fei Sun, Jiajun Tan, Xinyu Ma, Qi Cao, Dawei Yin, Huawei Shen, and Xueqi Cheng. 2025. The Mirage of Model Editing: Revisiting Evaluation in the Wild. CoRR abs/2502.11177 (2025). doi:10.48550/ARXIV.2502.11177 arXiv:2502.11177
- Yao et al. (2024) Ruoyu Yao, Yubin Wang, Haichao Liu, Rui Yang, Zengqi Peng, Lei Zhu, and Jun Ma. 2024. CALMM-Drive: Confidence-Aware Autonomous Driving with Large Multimodal Model. CoRR abs/2412.04209 (2024). doi:10.48550/ARXIV.2412.04209 arXiv:2412.04209
- Yao et al. (2025) Yunzhi Yao, Jizhan Fang, Jia-Chen Gu, Ningyu Zhang, Shumin Deng, Huajun Chen, and Nanyun Peng. 2025. CaKE: Circuit-aware Editing Enables Generalizable Knowledge Learners. arXiv preprint arXiv:2503.16356 (2025).
- Yao et al. (2023) Yunzhi Yao, Peng Wang, Bozhong Tian, Siyuan Cheng, Zhoubo Li, Shumin Deng, Huajun Chen, and Ningyu Zhang. 2023. Editing Large Language Models: Problems, Methods, and Opportunities. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 10222–10240. doi:10.18653/V1/2023.EMNLP-MAIN.632
- Youssef et al. (2025) Paul Youssef, Zhixue Zhao, Daniel Braun, Jörg Schlötterer, and Christin Seifert. 2025. Position: Editing Large Language Models Poses Serious Safety Risks. arXiv preprint arXiv:2502.02958 (2025).
- Zhang et al. (2024c) Junzhe Zhang, Huixuan Zhang, Xunjian Yin, Baizhou Huang, Xu Zhang, Xinyu Hu, and Xiaojun Wan. 2024c. MC-MKE: A Fine-Grained Multimodal Knowledge Editing Benchmark Emphasizing Modality Consistency. CoRR abs/2406.13219 (2024). doi:10.48550/ARXIV.2406.13219 arXiv:2406.13219
- Zhang et al. (2024a) Ningyu Zhang, Bozhong Tian, Siyuan Cheng, Xiaozhuan Liang, Yi Hu, Kouying Xue, Yanjie Gou, Xi Chen, and Huajun Chen. 2024a. InstructEdit: Instruction-Based Knowledge Editing for Large Language Models. In Proceedings of the Thirty-Third International Joint Conference on Artificial Intelligence, IJCAI 2024, Jeju, South Korea, August 3-9, 2024. ijcai.org, 6633–6641. https://www.ijcai.org/proceedings/2024/733
- Zhang et al. (2024b) Ningyu Zhang, Yunzhi Yao, Bozhong Tian, Peng Wang, Shumin Deng, Mengru Wang, Zekun Xi, Shengyu Mao, Jintian Zhang, Yuansheng Ni, Siyuan Cheng, Ziwen Xu, Xin Xu, Jia-Chen Gu, Yong Jiang, Pengjun Xie, Fei Huang, Lei Liang, Zhiqiang Zhang, Xiaowei Zhu, Jun Zhou, and Huajun Chen. 2024b. A Comprehensive Study of Knowledge Editing for Large Language Models. CoRR abs/2401.01286 (2024). doi:10.48550/ARXIV.2401.01286 arXiv:2401.01286
- Zhang et al. (2023) Qingru Zhang, Minshuo Chen, Alexander Bukharin, Pengcheng He, Yu Cheng, Weizhu Chen, and Tuo Zhao. 2023. Adaptive Budget Allocation for Parameter-Efficient Fine-Tuning. In The Eleventh International Conference on Learning Representations. https://openreview.net/forum?id=lq62uWRJjiY
- Zheng et al. (2023) Ce Zheng, Lei Li, Qingxiu Dong, Yuxuan Fan, Zhiyong Wu, Jingjing Xu, and Baobao Chang. 2023. Can We Edit Factual Knowledge by In-Context Learning?. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, 4862–4876. doi:10.18653/V1/2023.EMNLP-MAIN.296